OpenStudy (anonymous):

Sally estimated that her call would take 15 minutes but it actually took 12 minutes. What was her percent of error

OpenStudy (anonymous):

20 percent error 12/15=4/5=0.8 20 percent off

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Thanks Unimath and 14yamaka

OpenStudy (anonymous):

WELCOME XD

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

12/15 or she got 12 minutes right out of 15 giving .8, converting .8 to percent x100 equals 80 percent right leaving 20% error

OpenStudy (anonymous):

so many of the same post unimath0.0 what happened XD